# IndoCulture: Exploring Geographically-Influenced Cultural Commonsense Reasoning Across Eleven Indonesian Provinces

### v1

@@ Abstract
Although commonsense reasoning is greatly shaped by cultural and geographical factors, previous studies on language models have predominantly centered on English cultures, potentially resulting in an Anglocentric bias. In this paper, we introduce IndoCulture, aimed at understanding the influence of geographical factors on language model reasoning ability, with a specific emphasis on the diverse cultures found within eleven Indonesian provinces. In contrast to prior works that relied on template Yin et al. (2022) and online scrapping Fung et al. (2024), we create IndoCulture by asking local people to manually develop the context and plausible options based on predefined topics. Evaluations of 23 language models reveal several insights: (1) even the best open-source model struggles with an accuracy of 53.2%, (2) models often provide more accurate predictions for specific provinces, such as Bali and West Java, and (3) the inclusion of location contexts enhances performance, especially in larger models like GPT-4, emphasizing the significance of geographical context in commonsense reasoning. [^1] IndoCulture can be accessed at https://huggingface.co/datasets/indolem/IndoCulture
@@ 1 Introduction
Figure 1: IndoCulture covers eleven provinces spanning from eastern to western Indonesia. The highlighted regions in the map represent the provinces examined in IndoCulture. We offer examples from Aceh, South Borneo, and Papua, with correct answers indicated in bold. English translations are for illustrative purposes.
@@
Reasoning abilities of multilingual language models are frequently evaluated using English texts, potentially amplifying an Anglocentric bias toward English culture and leading to less inclusive models Thomas (1983); Ponti et al. (2020). Cultures, however, vary significantly from one location to another and profoundly shape the way a speaker of a language reasons Hershcovich et al. (2022). Recent evaluations of models reasoning ability OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Sengupta et al. (2023); Liu et al. (2023b) conducted on the English datasets, such as PIQA Bisk et al. (2020a), HellaSwag Zellers et al. (2019), Social IQA Sap et al. (2019), often overlook geographical aspects, thereby risking cultural bias.
@@
Culture is a multifaceted concept encompassing the way of life Giddens and Sutton (2021), including our thoughts and actions Macionis John . It includes tangible elements like food art, and clothing, as well as intangible aspects such as ideas, values, attitudes, and norms. Culture is shaped by geographical location and ethnicity, influencing the commonsense reasoning of people within a region. For example, in Indonesia, it is culturally acceptable to eat rice with your hands but it is considered unusual to use a fork. Similarly, at traditional Indonesian weddings, it is common to sit on the floor while eating, whereas this practice is less common in Australia.
@@
This work focuses on understanding the influence of geographical contexts in cultural commonsense reasoning, with the main focus on Indonesian culture. Indonesia is a highly multicultural country Putra et al. (2019), home to over 1,300 recognized ethnic groups and more than 700 languages Zarbaliyev (2017); Aji et al. (2022). As the the largest archipelagic country in the world, Indonesia has a population exceeding 270 million spread across 38 provinces, stretching from Aceh province in the west to Papua province in the east. Few prior studies on commonsense reasoning in Indonesian contexts Mahendra et al. (2021); Wibowo et al. (2023); Putri et al. (2024) overlooked the geographical nuance and the rich diversity of Indonesian cultures.
@@
This paper introduces IndoCulture, a novel dataset to evaluate cultural reasoning in eleven Indonesian provinces, manually developed by local people in each province based on predefined topics. In prior works, cultural reasoning has primarily relied on datasets constructed through templates Yin et al. (2022), and online scraping Nguyen et al. (2023a); Fung et al. (2024). While these studies offer valuable insights, they may be susceptible to training data contamination when utilized to assess large language models (LLMs). For instance, Fung et al. (2024) reported a zero-shot accuracy of 92% when using ChatGPT to evaluate low-resource data.
@@
Figure 1 illustrates the eleven provinces examined in IndoCulture including Aceh, North Sumatra, West Sumatra, West Java, Central Java, East Java, Bali, South Borneo, East Nusa Tenggara (NTT), South Sulawesi, and Papua. [^2] Although Papua consists of six provinces, for the purpose of this study, we treat it as a single entity (referred to as Papua) due to the relatively recent establishment of most of these provinces. These provinces span from the eastern to the western parts of Indonesia, each representing a major island in the country, with the addition of Bali and NTT. Figure 1 also shows three examples in IndoCulture for three provinces: Aceh, South Borneo, and Papua. The first two examples focus on cultural artifacts: the traditional wedding dress from Aceh and the Bagandut dance from South Borneo. The second example also examines cultural symbols. The third example focuses on cultural beliefs and norms regarding pregnancy in Papua.
@@
Can large language models effectively reason based on the diverse cultures of Indonesia? To capture the rich diversity of Indonesian cultures, we meticulously predefined 12 fine-grained topics as guidelines for data construction. Figure 2 displays the topic distribution in IndoCulture, with the majority focusing on food, weddings, art, pregnancy and children, and family relationships. Additionally, we also pose the question: Is there an influence of geographical location on commonsense reasoning by language models? We address these questions through comprehensive experiments across different language models, incorporating several levels of location granularity as additional context in the prompt.
@@
  • We present IndoCulture, a high-quality cultural reasoning dataset in the Indonesian language, covering eleven provinces of Indonesia and twelve fine-grained cultural topics. Our dataset has 2,429 instances, and was meticulously developed by local people with rigorous quality controls in place.
  • We assess 15 open-source multilingual models, 6 open-source Indonesian-centric models, and 2 closed-source models. Although local individuals can answer all questions correctly (i.e., 100% accuracy), we observed that all open-source models struggle to comprehend Indonesian cultures, with the highest performance reaching only 53.2%.
  • We conduct a thorough analysis over various dimensions: (1) models performance for each province and topics, (2) the influence of different granularities of location contexts (i.e., none, province, country), (3) model performance over English translation, (4) analysis over model explanation behind its answer.
@@ 2 Related Work
Many studies have focused on commonsense reasoning in English, often overlooking considerations of cultures and geographical locations. Early works include the Winograd Schema Challenge Levesque et al. (2012) and WinoGrande Sakaguchi et al. (2021) for pronoun coreference resolution tasks. Other research areas include reasoning based on cause-effect relationships Roemmele et al. (2011), physical activities Bisk et al. (2020b), social interaction Sap et al. (2019), cloze story completion Mostafazadeh et al. (2016), sentence completion Zellers et al. (2019), numerical Lin et al. (2020), and temporal Qin et al. (2021). Additionally, pretrained language models are employed in other works to extract structured commonsense knowledge by providing seed words Davison et al. (2019), and using code language model Madaan et al. (2022).
@@
Previous studies have explored commonsense reasoning with geographical contexts. Shwartz (2022) investigated time perception (e.g., morning and night hours) across different locations, while Yin et al. (2022) delved into cultural knowledge of language models across five countries by masking and template strategies. Other works have focused on automatically extracting cultural knowledge from various sources, including Wikipedia Fung et al. (2024), conversations Fung et al. (2023), and Common Crawl Nguyen et al. (2023a), incorporating location contexts with the assistance of large language models. Related, Ziems et al. (2023) created a knowledge bank for situational norms, utilizing English-speaking Mechanical Turk annotators and incorporating country taxonomy. Unlike these works, IndoCulture specifically concentrates on cultural reasoning within Indonesian provinces, developed and validated manually by local people (experts). Compared to the automatic method and English-speaking crowdsource workers for data construction, IndoCulture arguably contains less noise, and is free from training data contamination of large language models (LLMs).
@@
Commonsense reasoning in Indonesian language models has been studied using translated English–Indonesian datasets, such as XCOPA Ponti et al. (2020) and XStoryCloze Lin et al. (2021). However, these datasets potentially introduce a cultural bias toward English culture. IndoCloze Koto et al. (2022) is the first commonsense reasoning dataset in Indonesian, developed by native Indonesian workers following cloze story completion framework Mostafazadeh et al. (2016). IndoMMLU Koto et al. (2023), a benchmark for assessing knowledge in the Indonesian context, includes questions pertaining to Indonesian culture. However, its primary focus lies on testing proficiency in the local language rather than reasoning abilities. In a contemporary study, Wibowo et al. (2023) followed the COPA framework Roemmele et al. (2011) to build a dataset with contexts limited to Jakarta. In other work, Putri et al. (2024) studied the capability of LLMs in generating questions with cultural norms, for both general Indonesian and specific Sundanese contexts. In comparison, IndoCulture offers broader coverage across eleven provinces, thereby fostering greater inclusivity for local communities in Indonesia.
@@ 3 IndoCulture
As illustrated in Figure 1, IndoCulture is a sentence completion task in Indonesian language featuring a one-sentence premise, three plausible options, and one correct option to evaluate reasoning ability and cultural knowledge across eleven Indonesian provinces. In total, there are 2,429 instances, with the distribution across provinces detailed in Table 1.
@@ 3.1 Data Construction
IndoCulture is constructed manually by human and verified through a two-step process.
@@
Culture often arises from the shared experiences, traditions, and beliefs of a specific group over time, often closely intertwined with native populations. Keeping this in mind, we engaged individuals from various provinces across Indonesia to assist in preparing data for the IndoCulture benchmark. During the recruitment, we presented a few examples of the intended IndoCulture data and requested each candidate to generate similar instances tailored to the context of their respective provinces. From a pool of 58 applicants, we carefully selected 22 expert workers representing 11 provinces (with 2 workers selected per province). These recruited expert workers are local residents and have resided in their respective provinces for a minimum of 10 years, thereby possessing a profound understanding of local customs and culture. The age range of our workforce spans from 21 to 35 years old, with educational backgrounds distributed as follows: 3 high school graduates, 14 bachelor’s degree holders, 4 master’s degree holders, and 1 PhD holder. During data construction, each expert worker will fulfill the dual roles of instance writer and quality controller. [^3] Each worker is compensated above the monthly minimum wage in Indonesia.
@@
The provinces covered in this study represent diversity of Indonesian cultures. Those 11 provinces (in Figure 1) spread across 6 islands and large archipelagos in Indonesia which are inhabited by various tribes who speak different regional languages and adhere to different religions.
@@
IndoCulture consists of 12 topics and 66 fine-grained subtopics. Below is a list of topics along with their detailed subtopics. The numbers following the topic indicate the total instances required to be written by one worker (with a total of 150 per worker).
@@
    Food (22): breakfast (2); lunch (3); dinner (2); snacks (2); food souvenir (3); traditional foods and beverages (5); eating habit (1); cutlery (1); cooking ware (1), fruits (2).
@@
    Death (10): when death occurs (2); the process of caring for a corpse (2); traditions after the body is buried (2); the clothes of the mourners (2); inheritance matters (2).
@@
    Daily activities (10): morning activities (1); afternoon activities (1); evening activities (1); leisure activities (3); house, household and transportation (4).
@@
    Socio-religious aspects of life (13): regular religious activities (2); mystical things (2); traditional ceremonies (1); lifestyle (3); self care (1); traditional medicine (3); traditional saying (1).
@@
For each instance, workers craft two culturally relevant sentences that align with the predefined subtopic. The first sentence serves as the premise context, while the last sentence acts as the correct answer. Subsequently, the annotator generates two additional plausible sentences as distractors by modifying cultural objects or activities from the correct sentence. Workers are given a period of two months to complete the task.
@@
In stage 1, we implement quality control by pairing two annotators from the same province. Each annotator is tasked with answering a set of questions prepared by the other annotator, and vice versa. During this phase, the annotator is presented with a premise sentence and three shuffled options. They are allowed to search for the answer from any source if they are unsure. Instances that are incorrectly answered by the second annotator are discarded, as we hypothesize that these instances may contain incorrect answers or possess a level of ambiguity. Additionally, annotators are required to identify whether the instance is province-specific (binary annotation: True/False), meaning it is uniquely relevant in their province and not in others.
In stage 2 of quality control, the first two authors of this paper perform post-editing on the data that has passed the first stage of quality control. We first focus on correcting the linguistic aspects of the text, which includes checking for spelling typos. Although the data is written in Indonesian, some annotators may use dialects or be influenced by the structure or style of regional languages. In these cases, we correct the text to adhere to correct Indonesian grammar.
To maintain the quality of IndoCulture, we rigorously filter instances that display (1) poor writing, as enhancing their quality substantially is difficult; (2) obvious answer options, which allow for easy guessing of the correct choice without understanding the cultural contexts; and (3) ambiguous contexts, where all options are equally valid as the correct answer. For example, in a topic about breakfast, the three options might include one traditional food alongside two other very commonly consumed foods in Indonesia.
Furthermore, we manually verify the province-specific annotation for each instance using the Google search engine. We annotate whether the instance pertains to national-level culture. If the example is specific to a province, it will be annotated as uncommon in national culture, and vice versa.

### v2

@@ Abstract
Although commonsense reasoning is greatly shaped by cultural and geographical factors, previous studies have predominantly centered on cultures grounded in the English language, potentially resulting in an Anglocentric bias. In this paper, we introduce IndoCulture, aimed at understanding the influence of geographical factors on language model reasoning ability, with a specific emphasis on the diverse cultures found within eleven Indonesian provinces. In contrast to prior work that has relied on templates Yin et al. (2022) and online scrapping Fung et al. (2024), we create IndoCulture by asking local people to manually develop a cultural context and plausible options, across a set of predefined topics. Evaluation of 27 language models reveals several insights: (1) the open-weight Llama–3 is competitive with GPT–4, while other open-weight models struggle, with accuracies below 50%; (2) there is a general pattern of models generally performing better for some provinces, such as Bali and West Java, and less well for others; and (3) the inclusion of location context enhances performance, especially for larger models like GPT–4, emphasizing the significance of geographical context in commonsense reasoning. [^1] IndoCulture is available from https://huggingface.co/datasets/indolem/IndoCulture
@@ 1 Introduction
Figure 1: IndoCulture covers eleven provinces spanning from eastern to western Indonesia. The highlighted regions in the map represent the provinces examined in IndoCulture. We present examples from Aceh, North Sumatra, and Papua, with three plausible options and correct answers indicated in bold. English translations are provided for illustrative purposes.
@@
The reasoning abilities of multilingual language models are frequently evaluated using English texts, potentially amplifying an Anglocentric bias toward culture grounded in the English language, and leading to less inclusive models Thomas (1983); Ponti et al. (2020). Cultures, however, vary significantly from one location to another and profoundly shape the way speakers of a language reason Hershcovich et al. (2022). Recent evaluations of models’ commonsense reasoning ability OpenAI (2023); Sengupta et al. (2023); Liu et al. (2023) have been conducted on English datasets such as Social IQA Sap et al. (2019) and PIQA Bisk et al. (2020), and thus often overlook geographical aspects, thereby risking cultural bias.
@@
Culture is a multifaceted concept encompassing the way of life Giddens and Sutton (2021), including our thoughts and actions Macionis (2012). It includes tangible elements like food, art, and clothing, as well as intangible aspects such as ideas, values, attitudes, and norms. Culture is shaped by geographical location and ethnicity, influencing the commonsense reasoning of people within a region. For example, in Indonesia, it is culturally acceptable to eat rice with your hands but it is considered unusual to use chopsticks. Similarly, at traditional Indonesian weddings, it is common to sit on the floor while eating, whereas this practice is less common in Australia.
@@
This work focuses on understanding the influence of geographical contexts in cultural commonsense reasoning, with the main focus on Indonesian culture. Indonesia is a highly multicultural country Putra et al. (2019), home to over 1,300 recognized ethnic groups and more than 700 languages Zarbaliyev (2017); Aji et al. (2022). As the the largest archipelagic country in the world, Indonesia has a population exceeding 270 million spread across 38 provinces, stretching from Aceh province in the west to Papua province in the east. Few prior studies on commonsense reasoning in Indonesian contexts Mahendra et al. (2021); Wibowo et al. (2024); Putri et al. (2024) have explicitly addressed the geographical nuances and rich diversity of Indonesian cultures.
@@
This paper introduces IndoCulture, a novel dataset to evaluate cultural reasoning in eleven Indonesian provinces, manually developed by local people in each province based on predefined topics. In prior work, cultural reasoning has primarily relied on datasets constructed through templates Yin et al. (2022), and online scraping Nguyen et al. (2023); Fung et al. (2024). While these studies offer valuable insights, they may be susceptible to training data contamination when used to assess large language models (LLMs). For instance, Fung et al. (2024) reported a zero-shot accuracy of 92% when using ChatGPT Ouyang et al. (2022) to evaluate low-resource data.
@@
IndoCulture contains cultural commonsense knowledge data from eleven provinces in Indonesia (blue colored in Figure 1), namely Aceh, North Sumatra, West Sumatra, West Java, Central Java, East Java, Bali, South Borneo, East Nusa Tenggara (NTT), South Sulawesi, and Papua. These provinces span breadth of Indonesia, each representing a major island in the country, with the addition of Bali and NTT. Figure 1 also shows three examples in IndoCulture for three provinces: Aceh, North Sumatra, and Papua. [^2] Although Papua consists of six provinces, for the purpose of this study, we treat it as a single entity (referred to as Papua) due to the relatively recent establishment of most of these provinces. The first example focuses on cultural artifact, specifically the traditional wedding dress from Aceh. The second example examines family relationships while the third example focuses on cultural beliefs and norms regarding pregnancy in Papua.
Dataset Size Data Construction Method Cultural? Location? #province #topic
IndoCulture (ours) 2,429 Manually built and validated by native 11 66
COPAL-ID Wibowo et al. (2024) 559 Manually built and validated by native
MAPS Liu et al. (2024) 371 LLM generation & human generation 1
ID-CSQA Putri et al. (2024)* 4,416 LLM generation & human generation 5
BLEnD Myung et al. (2024) 1,000 Template, translation, human validation 1 6
IndoCloze Koto et al. (2022) 2,335 Manually built and validated by native
XCOPA Ponti et al. (2020) 600 Translated from English data
XStoryCloze Lin et al. (2022) 1,872 Translated from English data
Table 1: Comparison of IndoCulture with other cultural knowledge and reasoning datasets containing instances in Indonesian. The metadata includes Size (number of Indonesian instances), Cultural? (whether the data considers cultural nuances), Location? (whether the data includes fine-grained location information, such as provinces, as context), #province (number of Indonesian provinces covered), and #topic (number of fine-grained topics covered). * indicates the dataset involves question generation with less emphasis on reasoning.
@@
Can large language models effectively reason based on the diverse cultures of Indonesia? To capture the rich diversity of Indonesian cultures, we predefined 12 fine-grained topics as guidelines for data construction. Figure 2 displays the topic distribution in IndoCulture, with the majority focusing on food, weddings, art, pregnancy and children, and family relationships. Additionally, we also pose the question: Is there any influence of geographical location on the commonsense reasoning of language models? We address these questions through comprehensive experiments across different language models, incorporating several levels of location granularity as additional context in the prompt.
@@
  • We present IndoCulture, a high-quality cultural reasoning dataset in the Indonesian language, covering eleven provinces of Indonesia and twelve fine-grained cultural topics. Our dataset has 2,429 instances, and was developed by local people with rigorous quality controls in place.
  • We assess 19 open-weight multilingual models, 6 open-weight Indonesian-centric models, and 2 closed-weight models. Although local individuals can answer all questions correctly (i.e., 100% accuracy), most open-weight models struggle to comprehend Indonesian cultures. Interestingly, we observed that Llama–3 Dubey et al. (2024) is competitive with GPT–4 OpenAI (2023).
  • We conduct a thorough analysis over various dimensions: (1) model performance for each province and topic; (2) the influence of different granularities of location context (i.e., none, province, country); (3) model performance over English translations; and (4) analysis of model explanations for a given answer.
@@ 2 Related Work
Many studies have focused on commonsense reasoning in English, often overlooking considerations of culture and geographical location. Early work included the Winograd Schema Challenge Levesque et al. (2012) and WinoGrande Sakaguchi et al. (2021) for pronoun coreference resolution. Other research areas include reasoning based on cause-effect relationships Roemmele et al. (2011), physical activities Bisk et al. (2020), social interactions Sap et al. (2019), cloze story completion Mostafazadeh et al. (2016), sentence completion Zellers et al. (2019), numerical reasoning Lin et al. (2020), and temporal reasoning Qin et al. (2021). Additionally, pretrained language models have been employed in other work to extract structured commonsense knowledge by providing seed words Davison et al. (2019), and using code language models Madaan et al. (2022).
@@
Previous studies have explored commonsense reasoning with geographical context. Shwartz (2022) investigated time perception (e.g., morning and night) across different locations, while Yin et al. (2022) examined cultural knowledge of language models across five countries using datasets built from templates and translations. Other work has focused on automatically extracting cultural knowledge from various sources, including Wikipedia Fung et al. (2024), conversations Fung et al. (2023), and Common Crawl Nguyen et al. (2023), incorporating location context with the assistance of large language models. Related, Ziems et al. (2023) created a knowledge bank for situational norms, using English-speaking Mechanical Turk annotators and incorporating a country taxonomy. Unlike this work, IndoCulture specifically concentrates on cultural reasoning across Indonesian provinces, developed and validated manually by local people (experts). Compared to the automatic method and English-speaking crowd workers for data construction, IndoCulture arguably contains less noise, and is free from the training data contamination of large language models (LLMs).
@@
Table 1 shows a comparison of IndoCulture with other Indonesian datasets for cultural knowledge and reasoning evaluation. Commonsense reasoning in Indonesian language models has been studied using translated English–Indonesian datasets, such as XCOPA Ponti et al. (2020) and XStoryCloze Lin et al. (2022). However, these datasets potentially introduce a cultural bias toward culture grounded in the English language. IndoCloze Koto et al. (2022) was the first commonsense reasoning dataset in Indonesian, developed by native Indonesian workers following the cloze story completion framework Mostafazadeh et al. (2016). However, IndoCloze lacks local cultural nuances and fine-grained geographical context. Wibowo et al. (2024) followed the COPA framework Roemmele et al. (2011) to build a dataset with contexts limited to Jakarta. In other work, Putri et al. (2024) studied the capability of LLMs in generating questions with cultural norms, for both general Indonesian and specific Sundanese contexts, while Liu et al. (2024) used proverbs and LLMs to generate conversational data. In contemporary work, Myung et al. (2024) released BLEnD, a large-scale cultural knowledge dataset, built using templates, translation, and human validations, covering the West Java province in Indonesia. BLEnD specifically focuses on short-answer questions, limiting its capacity for reasoning evaluation. Unlike most other datasets that do not consider geographical factors, IndoCulture has broad coverage across eleven provinces, thereby providing greater inclusivity for local communities in Indonesia.
@@ 3 IndoCulture
As illustrated in Figure 1, IndoCulture is a sentence completion task in the Indonesian language featuring a one-sentence premise, three plausible options, and one correct option to evaluate reasoning ability and cultural knowledge across eleven Indonesian provinces. While sentence completion tasks are straightforward for humans, answering IndoCulture requires machines to engage in cultural reasoning to logically conclude which of the three options is logically consistent with the first sentence Huang and Chang (2023). The dataset includes a total of 2,429 instances.
@@ 3.1 Data Construction
IndoCulture was constructed manually by humans, and verified through a two-step process.
@@
Culture generally arises from the shared experiences, traditions, and beliefs of a specific group over time, often closely intertwined with native populations. With this in mind, we engaged individuals from various provinces across Indonesia to assist in preparing data for the IndoCulture benchmark.
During recruitment, we presented a few examples of the intended IndoCulture data and requested each candidate to generate similar instances tailored to the context of their respective provinces. From a pool of 58 applicants, we carefully selected 22 expert workers representing 11 provinces (with 2 workers selected per province). These recruited expert workers are local residents and have resided in their respective provinces for a minimum of 10 years, thereby possessing a profound understanding of local customs and culture. The age range of our workforce spans from 21 to 35 years old, with educational backgrounds distributed as follows: 3 high school graduates, 14 bachelor’s degree holders, 4 Master’s degree holders, and 1 PhD holder.
During data construction, each expert worker fulfilled the dual roles of instance writer and quality controller. Each worker was compensated above the monthly minimum wage in Indonesia.
@@
The provinces covered in this study represent the diversity of Indonesian cultures. The 11 provinces (in Figure 1) are spread across 6 islands of the Indonesian archipelago, which are inhabited by different ethnic groups who speak different regional languages and adhere to different religions.
@@
IndoCulture consists of 12 topics and 66 fine-grained subtopics, carefully constructed based on discussions and brainstorming with Indonesian natives. The selection of these topics and subtopics was guided based on several criteria and motivations: (1) relevance to Indonesian culture; (2) diversity and coverage; (3) regional representation (e.g., religious holidays); (4) practicality; and (5) expert consultation (i.e., native speaker feedback). Compared to the other Indonesian datasets in Table 1, IndoCulture includes a richer array of fine-grained topics. Below is a list of the topics along with their detailed subtopics. The numbers following each topic indicate the total number of instances required to be written by one worker (with a total of 150 per worker).
@@
    Food (22): breakfast (2); lunch (3); dinner (2); snacks (2); food souvenirs (3); traditional foods and beverages (5); eating habits (1); cutlery (1); cooking ware (1), fruit (2).
@@
    Death (10): when death occurs (2); the process of dealing with a corpse (2); traditions after the body is buried (2); the clothes of the mourners (2); inheritance matters (2).
@@
    Daily activities (10): morning activities (1); afternoon activities (1); evening activities (1); leisure activities (3); house, household, and transportation (4).
@@
    Socio-religious aspects of life (13): regular religious activities (2); mystical things (2); traditional ceremonies (1); lifestyle (3); self care (1); traditional medicine (3); traditional sayings (1).
@@
For each instance, workers were asked to craft two culturally-relevant sentences that align with the predefined subtopic. The first sentence serves as the premise context, while the last sentence acts as the correct answer. Subsequently, the annotator generates two additional plausible sentences as distractors by modifying cultural objects or activities from the correct sentence. These distractors are designed to reflect local cultural contexts, ensuring they are challenging yet unambiguous, and could potentially serve as correct answers in other regional contexts. Workers were given a period of two months to complete the task.
@@
In stage 1, we implemented quality control by pairing two annotators from the same province. Each annotator was tasked with answering a set of questions prepared by the other annotator, and vice versa. During this phase, the annotator were presented with a premise sentence and three shuffled options. They were allowed to search for the answer from any source if they were unsure. Instances that were incorrectly answered by the second annotator were discarded, as we hypothesize that these instances may contain incorrect answers or possess a level of ambiguity. Additionally, annotators were required to identify whether the instance is province-specific (binary annotation: True/False), indicating that it is uniquely relevant in their province and not in others.
In stage 2 of quality control, the first two authors of this paper performed post-editing of data that passed the first stage of quality control. We first focused on correcting the linguistic aspects of the text, including checking for spelling errors. Although the text is written in Indonesian, some annotators may use dialects or be influenced by the structure or style of regional languages. In these cases, we corrected the text to adhere to Indonesian grammar.
To maintain the quality of IndoCulture, we rigorously filtered instances that contained: (1) poor writing, in the case that it was difficult to post-edit to enhance their quality; (2) obvious answer options, which allow for easy guessing of the correct choice without understanding the cultural context; and (3) ambiguous contexts, where all options are equally valid as the correct answer. For example, in a topic about breakfast, the three options might include one traditional food alongside two other very commonly consumed foods in Indonesia, and be considered too obvious.
Furthermore, we manually verified the province-specific annotations for each instance using the Google search engine. We annotated whether the instance pertains to national-level culture or not. If the example is specific to a province, it will be annotated as uncommon in national culture, and vice versa.